# Policy Learning with a Language Bottleneck

### v2

@@ Abstract
Modern AI systems such as self-driving cars and game-playing agents achieve superhuman performance, but often lack human-like generalization, interpretability, and inter-operability with human users. Inspired by the rich interactions between language and decision-making in humans, we introduce Policy Learning with a Language Bottleneck (PLLB), a framework enabling AI agents to generate linguistic rules that capture the high-level strategies underlying rewarding behaviors. PLLB alternates between a rule generation step guided by language models, and an update step where agents learn new policies guided by rules, even when a rule is insufficient to describe an entire complex policy. Across five diverse tasks, including a two-player signaling game, maze navigation, image reconstruction, and robot grasp planning, we show that PLLB agents are not only able to learn more interpretable and generalizable behaviors, but can also share the learned rules with human users, enabling more effective human-AI coordination. [^1] We provide source code for our experiments at https://github.com/meghabyte/bottleneck.
@@ 1 Introduction
As AI systems play an increasingly central role in automation, there is a growing need for them to better model and emulate human behavior. To be truly effective, these systems must be capable of acting and generalizing in ways that align with human expectations. Unfortunately, many of today’s AI systems do not meet these standards. Self-driving cars or game-playing agents may achieve super-human performance, but they lack interpretability (McIlroy-Young et al., 2020) and often act unpredictably, especially outside their training distribution (Wang et al., 2023).
Figure 1: Policy Learning with a Language Bottleneck (PLLB) alternates between two steps: 1) gen_rule (bottom) generates a linguistic rule $\mathcal{L}_{i}$ explaining the agent’s best behaviors by prompting a language model with contrastive (positive and negative) episodes (middle); 2) update (top) learns a new policy conditioned on that rule. PLLB facilitates stronger human-AI coordination by constraining policies to be more interpretable and generalize better.
Humans acquire most of their skills and knowledge from others, often using language — via instructions, advice, or explanations that improve their decision-making capabilities (Carruthers and Boucher, 1998; Mesoudi and Thornton, 2018). Language acts as a communicative medium, enabling us to teach, learn from, and coordinate with others to solve complex problems. It also supports other cognitive functions: even when not used for communication, it allows us to represent abstract concepts (Hesse, 1988; Lakoff and Johnson, 2008), and plan (Vygotsky, 1965; Clark, 1998); it guides our attention (Waxman, 1994; Yoshida and Smith, 2003), and prompts relational thinking (Gentner and Loewenstein, 2002).
@@
Recent literature contains many examples of AI systems leveraging language-based representations or linguistic feedback, but these often rely on external supervision: humans in the loop or hard-coded feedback functions (Luketina et al., 2019; Colas et al., 2022). We argue that more human-like AI systems should not only use language-based supervision but also generate their own language-based feedback to leverage both the communicative and cognitive functions of language. This capability should also extend to tasks that require updating an underlying policy that is only partially expressible in natural language.
This paper introduces Policy Learning with a Language Bottleneck (PLLB), a framework that providesides artificial embodied agents the ability to generate linguistic rules that capture the strategies underlying their most rewarding behaviors. As shown in Figure 1, PLLB alternates between a rule generation step that explains the agent’s experiences by prompting a language model (LM) with contrastive episodes, and a policy update step that learns a new policy guided by these rules. Unlike past work that solely leverages LMs for modeling agent behavior and multistep reasoning (Park et al., 2023; Wei et al., 2023), PLLB is applicable even when aspects of the target policy cannot be expressed with language.
Policy Learning with a Language Bottleneck can be applied to a wide range of agent types, from RL policies to LLM-based learners to robot pose estimators. We thus investigate its role in shaping more human-like policies across five distinct tasks. They perform better: in two image reconstruction tasks, PLLB agents generate instructions increasing the listeners’ performance compared to non-linguistic baselines (Section 7), and PLLB agents also help more efficiently learn robot grasping policies (Section 8). They generalize better: in a maze task, PLLB helps uncover the abstract problem structure and generalize better to similar mazes (Section 6), while also reducing overreliance on non-generalizable visual features in robotic manipulation (Section 8). They are more interpretable: in a coordination task, agents converge on humans’ preferred policy when multiple optimal policies exist (Section 5). They are more inter-operable: in maze and image reconstruction tasks, humans achieve better rewards when interacting with PLLB agents compared to agents trained without a bottleneck (Sections 6 and 7).
@@ 2 Background & Related Work
Multi-step reasoning and text agents. Recent works have proposed guiding LMs’ reasoning by prompting them to step through sequences of “thoughts” (Wei et al., 2023; Yao et al., 2022; Li et al., 2023b; Shinn et al., 2024). Our approach similarly uses a language bottleneck to concisely express intermediate information useful for later behavior (i.e. policy learning). However, unlike text-based reasoning agents such as ReAct (Yao et al., 2022), Rememberer (Zhang et al., 2023), and Reflexion (Shinn et al., 2024), PLLB does not require the underlying policy to be fully expressible in text. For example, we show that a text-only maze solving agent completely fails, whereas PLLB can improve classical Q-learning style agents (Section 6). Moreover, instead of describing actions for a particular state, PLLB rules efficiently capture high-level strategies over a sequence of states and actions and can generalize in similar settings. We show the learned rules can be shared to improve human task performance across multiple user studies (Sections 6 and 7).
@@ 3 The Language Bottleneck
Figure 2: PLLB can be applied to a diversity of domains. In all domains, it iterates between gen_rule, a function that prompts an LM to extract a linguistic rule ($\mathcal{L}$, blue) by contrasting high and low reward episodes from past experience (green and red boxes), and update, a function for updating an agent’s policy through interaction with the environment conditioned on $\mathcal{L}$. PLLB can be applied to multi-step decision making tasks and visual or robotics tasks alike with minimal implementation variations (replacing the LM by a VLM in gen_rule, or replacing policy regularization by simple instruction conditioning in update. Exact prompts are in Appendix Section A.5.
@@
PLLB builds on the standard RL framework to train agents to solve decision-making tasks in human-like ways. We formalize decision-making tasks (e.g.,  solving a maze) as Markov decision processes $(S,A,f,R,T)$ with reward function $R:S\times A\rightarrow\mathbb{R}$ (e.g.,  solving speed) over states $S$ (e.g.,  cell coordinates) and actions $A$ (e.g.,  directions), finite time horizon $T$, and a deterministic transition function $f:S\times A\rightarrow S$ that maps state–action pairs $(s,a)$ to next states $s$ (e.g.,  moving between cells). Standard RL then involves training a policy $\pi$ to maximize expected reward (e.g.,  learning to solve the maze faster). This is usually done by alternating between two steps: (1) collecting data with the current policy: $D\leftarrow\pi_{i}$ and (2) updating the policy using the data: $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D)$, where update implements some RL algorithm (Sutton and Barto, 2018). This procedure often fails to yield interpretable, human inter-operable, or generalizable behaviors.
@@
Our key idea is to introduce a language bottleneck between data collection and policy update. We extract linguistic rules explaining past rewarding behaviors and then use them to regularize the policy’s behavior in the next learning iteration (e.g., I should go right in blue cells, see Figure 1). Our experiments will show that this improves the interpretability and inter-operability (communicative use), and performance and generalization (cognitive use) of our agents over standard RL baselines in a variety of tasks (see Figure 2 and Sections 5 to 7). The resulting algorithm thus alternates between three steps: (1) data collection, $D_{i}\leftarrow\pi_{i}$; (2) language bottleneck generation, $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D_{i})$ and (3) policy updating, $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$, see Figure 1. In PLLB, the first step is implemented just as in ordinary policy learning approaches; we next describe the last two steps in more detail.
@@ 3.1 Rule Generation (gen_rule)
Using all the experience $D_{i}$ collected by the policy $\pi_{i}$ in the current iteration, gen_rule aims to infer an abstract rule $\mathcal{L}_{i}$ that best explains the agents’ successful behaviors: $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D)$. This is done by prompting an LM with contrastive episodes from $D_{i}$ (top-$N$ highest vs. top-$N$ lowest total rewards) and asking it to provide the rule that should be followed to obtain high rewards (see first row of Figure 2 and Appendix Section A.5 for full prompts). Importantly, this requires the first iteration of gen_rule to start only once we observe a pair episodes with sufficiently different rewards. We found this contrastive approach, inspired by Zhong et al. (2023) and Dunlap et al. (2023), to provide more precise rules than simply summarizing high-reward strategies.
@@ 3.2 Rule-Guided Policy Update (update)
Given a rule $\mathcal{L}_{i}$, the update step produces a new policy $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$ that is better aligned with $\mathcal{L}_{i}$. There exist many methods for leveraging language instructions, typically provided by experts, to modify agents’ policies. One approach we leverage is InstructRL (Hu and Sadigh, 2023), which regularizes the learned policy with another policy induced by the linguistic rule $\pi_{\mathcal{L}}$. For instance, in the maze example shown in Figure 1, if the rule is I should go right on every blue cell, the induced policy $\pi_{\mathcal{L}}$ should assign probability 1 to the right action in blue cells, but equal probabilities to all actions in every other situations. In the Q-learning algorithm, this approach simply adds a regularizing term (orange) to the standard Q-learning update rule  [^2] This regularizing term can also be applied in the function approximation setting (Mnih et al., 2013).:
@@
$Q^{\theta}(s_{t},a_{t})\leftarrow r_{\text{t+1}}+\gamma Q^{\theta}(s_{t+1},a_{t+1}),\textrm{ where}$
$a_{t+1}=\arg\max_{a}[Q(s_{\text{t+1}},a){\color[rgb]{1,.5,0}+\lambda\log\pi_{\mathcal{L}}(a\mid s_{t})}].$
@@
Here $\gamma$ is a discount factor and $\lambda$ controls the strength of the rule-induced regularization and could be made time-dependent with a pre-defined or learned schedule (e.g., stronger regularization at the beginning of learning).
@@
But how do we induce $\pi_{\mathcal{L}}$ from $\mathcal{L}$? Since $\mathcal{L}$ is expressed in natural language, we emulate it with an LM. We prompt the LM with both the rule and the current state of the agent $s_{t}$ and instruct it to generate the next action to obtain a probability distribution over admissible actions (e.g directions). This distribution can be constructed by normalizing the probability of generating each of the action tokens, as in InstructRL (Hu and Sadigh, 2023). Alternatively, for more complex action spaces, we can construct an estimate around the frequency of actions over repeated inference. Running Q-learning updates with the regularization term from experience data gives us a new Q-table $Q^{\theta}_{i+1}$ from which we can derive the new policy $\pi_{i+1}$ by taking actions with maximum expected value in every state $\pi_{i+1}(s_{t})=\text{max}_{a}Q^{\theta}_{t+1}(s_{t},a)$.
@@
Finally, there exist some domains where the policy can be directly implemented by an LM or VLM, which is leveraged by existing methods such as ReAct (Yao et al., 2022) or Rememberer (Zhang et al., 2023). In these cases, simply conditioning the policy on the rule $\mathcal{L}$ by adding it to the prompt $\pi_{\mathcal{L}}$ might be enough to steer the agent’s behavior to become more human-like: $\pi_{i+1}\leftarrow\pi_{\mathcal{L}_{i}}$. Given a new policy $\pi_{i+1}$, we can now close the loop and generate a new rule $\mathcal{L}_{i+1}$. We not only implement PLLB with such policies based on LM/VLMs in Section 7 , but also with classical RL agents (Sections 5 and 6), and robotics (Section 8).
@@ 4 Experiment Set-Up
We run experiments on a diverse set of five tasks, with different agent architectures and action spaces, to showcase how PLLB can train more human-like agents: a simple two-player communication game called SaySelect (Section 5), a Maze solving task (Section 6), two collaborative image reconstruction tasks with synthetic (Builder) and natural (Birds) images (Section 7), and grasp planning for a 7-DoF robot (Section 8). For all tasks, we include hyperparameter details (Appendix A.1), the exact prompts used in gen_rule and update (Appendix A.5), and examples of generated rules, including a qualitative analysis on the effect of the underlying language model on gen_rule (Appendix B.1 and B.2).
To evaluate the cognitive function of PLLB, we first evaluate the policies learned by artificial agents in each environment based on performance metrics such as task reward, similarity to known human-like policies, or generalization. For all tasks, we compare our Bottleneck approach with baseline methods (e.g., standard RL, base LM) and an Adversarial ablation of PLLB that leverages invalid linguistic rules (e.g., obtained by labeling high reward examples as low reward) in gen_rule.
To evaluate the communicative function of PLLB, we conduct several human subject studies showing that generated rules $\mathcal{L}$ can help improve human task performance on a variety of metrics (e.g., reward, solving time). Human subject study participants were either recruited in-person (for Maze) or on the Prolific crowdsourcing platform [^3] Prolific, http://prolific.com. (Builder, Birds), compensated at a rate of US\$12/hour. All studies were approved by our institution’s IRB.

### v3

@@ Abstract
Modern AI systems such as self-driving cars and game-playing agents achieve superhuman performance. But they often lack human-like generalization, interpretability, and inter-operability with human users. This paper introduces Policy Learning with a Language Bottleneck (PLLB), a framework enabling AI agents to generate linguistic rules that capture the high-level strategies underlying rewarding behaviors. PLLB alternates between a rule generation step guided by language models, and an update step where agents learn new policies guided by rules. Crucially, PLLB enables this kind of language-guided learning even when a natural language rule is insufficient to completely describe the target policy. Across five diverse tasks, including a two-player signaling game, maze navigation, image reconstruction, and robot grasp planning, we show that PLLB learns more interpretable and generalizable behaviors than standard policy learning methods. In three additional human subject studies, we show that show the learned rules significantly improve human task performance, enabling more effective human-AI coordination. [^1] We provide source code for our experiments at https://github.com/meghabyte/bottleneck.
Figure 1: Policy Learning with a Language Bottleneck (PLLB) alternates between two steps: 1) gen_rule generates a linguistic rule $\mathcal{L}_{i}$ explaining the agent’s best behaviors by prompting a language model with contrastive (positive and negative) episodes; 2) update learns a new policy conditioned on $\mathcal{L}_{i}$. PLLB strengthens human-AI coordination by constraining policies to be more interpretable and generalize better.
@@ 1 Introduction
As AI systems play an increasingly central role in automation and decision-making, their success depends not only on their performance, but also on their ability to model and align with human behavior. To be truly effective assistants, these systems must act and generalize in ways that align with human expectations. However, many of today’s AI systems do not meet these standards. Self-driving cars or game-playing agents like AlphaZero may achieve super-human performance, but they lack interpretability (McIlroy-Young et al., 2020) and often act unpredictably, especially outside their training distribution (Wang et al., 2023).
Unlike AI systems that are trained in isolation, humans acquire most of their skills and knowledge from interacting with others, often using language — via instructions, advice, or explanations that improve their decision-making capabilities (Carruthers and Boucher, 1998; Mesoudi and Thornton, 2018). Language acts as a communicative medium, enabling us to teach, learn from, and coordinate with others to solve complex problems. It also supports other cognitive functions: even when not used for communication, it allows us to represent abstract concepts (Hesse, 1988; Lakoff and Johnson, 2008), and plan (Vygotsky, 1965; Clark, 1998); it guides our attention (Waxman, 1994; Yoshida and Smith, 2003), and prompts relational thinking (Gentner and Loewenstein, 2002).
@@
There exist many recent examples in the literature of AI systems leveraging language-based representations or linguistic feedback, but these often rely on external supervision: humans in the loop or hard-coded feedback functions (Luketina et al., 2019; Colas et al., 2022). We argue that more human-like AI systems should not only use language-based supervision but also generate their own language-based feedback to leverage both the communicative and cognitive functions of language. Importantly, this capability should also extend to tasks that require updating an underlying policy that is only partially expressible in natural language, such as low-level control or reflexive actions.
This paper introduces Policy Learning with a Language Bottleneck (PLLB), a framework that provides artificial embodied agents the ability to generate linguistic rules that capture the strategies underlying their most rewarding behaviors. As shown in Figure 1, PLLB alternates between a rule generation step that explains the agent’s experiences by prompting a language model (LM) with contrastive episodes, and a policy update step that learns a new policy guided by these rules. Unlike past work that solely leverages LMs for modeling agent behavior and multistep reasoning (Park et al., 2023; Wei et al., 2023), PLLB is applicable even when aspects of the target policy cannot be expressed with language.
Policy Learning with a Language Bottleneck can be applied to a wide range of agent types, from RL policies to LLM-based learners to robot pose estimators, and the core mechanism of PLLB remains the same: using contrastive reward signals to extract linguistic rules via an LLM, and then conditioning policy updates on these rules depending on the type of learning agent. We investigate the role of PLLB in shaping more human-like policies across five distinct tasks. They perform better: in two image reconstruction tasks, PLLB agents generate instructions increasing the listeners’ performance compared to non-linguistic baselines (Section 7), and PLLB agents also help more efficiently learn robot grasping policies (Section 8). They improve few-shot generalization: in a maze task, PLLB rules uncover abstract problem structure that improve learning similar mazes (Section 6), and PLLB also reduces reliance on non-generalizable visual features in robotic manipulation (Section 8). They are more interpretable: in a coordination task, agents converge on humans’ preferred policy when multiple optimal policies exist (Section 5). They are more inter-operable: in maze and image reconstruction tasks, humans achieve better rewards when interacting with PLLB agents compared to agents trained without a bottleneck (Sections 6 and 7).
@@ 2 Background & Related Work
Multi-step reasoning and text agents. Recent works have proposed guiding LMs’ reasoning by prompting them to step through sequences of “thoughts” (Wei et al., 2023; Yao et al., 2022; Li et al., 2023b; Shinn et al., 2024). Our approach similarly uses a language bottleneck to concisely express intermediate information useful for later behavior (i.e. policy learning). However, unlike text-based reasoning agents such as ReAct (Yao et al., 2022), Rememberer (Zhang et al., 2023), and Reflexion (Shinn et al., 2024), PLLB does not require the underlying policy to be fully expressible in text. For example, we show that a text-only maze solving agent completely fails, whereas PLLB can improve classical Q-learning style agents (Section 6). Moreover, instead of describing actions for a particular state, PLLB rules capture high-level strategies over a sequence of states and actions, enabling generalization. We also show the learned rules can be shared to improve human task performance across multiple user studies (Sections 6 and 7).
@@ 3 The Language Bottleneck
Figure 2: PLLB can be applied to a diversity of domains. In all domains, it iterates between gen_rule, a function that prompts an LM (top gray boxes) to extract a linguistic rule ($\mathcal{L}$, blue) by contrasting high and low reward episodes from past experience (green and red boxes), and update, a function for updating an agent’s policy through interaction with the environment conditioned on $\mathcal{L}$. PLLB can be applied to multi-step decision making tasks and visual or robotics tasks alike with minimal implementation variations (replacing the LM by a VLM in gen_rule, or replacing policy regularization by simple instruction conditioning in update. Gray boxes represent prompts gen_rule and update, and full prompt details are in Appendix Section B.7.
@@
PLLB builds on the standard RL framework to train agents to solve decision-making tasks in human-like ways. We formalize decision-making tasks (e.g., solving a maze) as Markov decision processes $(S,A,f,R,T)$ with reward function $R:S\times A\rightarrow\mathbb{R}$ (e.g., solving speed) over states $S$ (e.g., cell coordinates) and actions $A$ (e.g., directions), finite time horizon $T$, and a deterministic transition function $f:S\times A\rightarrow S$ that maps state–action pairs $(s,a)$ to next states $s$ (e.g., moving between cells). Standard RL then involves training a policy $\pi$ to maximize expected reward (e.g., learning to solve the maze faster). This is usually done by alternating between two steps: (1) collecting data with the current policy: $D\leftarrow\pi_{i}$ and (2) updating the policy using the data: $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D)$, where update implements an RL algorithm (Sutton and Barto, 2018). This procedure often fails to yield interpretable, human inter-operable, or generalizable behaviors.
@@
Our key idea is to introduce a language bottleneck between data collection and policy update. We extract linguistic rules explaining past rewarding behaviors and then use them to regularize the policy’s behavior in the next learning iteration (e.g., I should go right in blue cells, see Figure 1). Our experiments will show that this improves the interpretability and inter-operability (communicative use), and performance and generalization (cognitive use) of our agents over standard RL baselines in a variety of tasks (see Figure 2 and Sections 5 to 7). The resulting algorithm thus alternates between three steps: (1) data collection, $D_{i}\leftarrow\pi_{i}$, which is implemented just as in ordinary policy learning approaches; (2) language bottleneck generation, $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D_{i})$ and (3) policy updating, $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$ (see Figure 1).
The core mechanism underlying PLLB is consistent across all instantiations: gen_rule prompts a language model with contrastive examples (high- vs. low-reward trajectories) to extract a rule $\mathcal{L}$ capturing what distinguishes successful from unsuccessful behavior, and update conditions the agent’s subsequent learning on $\mathcal{L}$. What varies across domains is only how update incorporates the rule, which is depends on the learning agent’s modality in that domain (e.g. Q-value regularization (Sections 56), prompt conditioning for LMs (Section 7) or visuomotor control (Section 8). We next describe these the gen_rule and update steps in more detail.
@@ 3.1 Rule Generation (gen_rule)
Using all the experience $D_{i}$ collected by the policy $\pi_{i}$ in the current iteration, gen_rule aims to infer an abstract rule $\mathcal{L}_{i}$ that best explains the agents’ successful behaviors: $\mathcal{L}_{i}\leftarrow\texttt{gen\_rule}(D)$. This is done by prompting an LM with contrastive episodes from $D_{i}$ (top-NN highest vs. top-$N$ lowest total rewards) and asking it to provide the rule that should be followed to obtain high rewards (see first row of Figure 2 and Appendix Section B.7 for full prompts). Importantly, this requires the first iteration of gen_rule to start only once we observe a pair episodes with sufficiently different rewards. We found this contrastive approach, inspired by Zhong et al. (2023) and Dunlap et al. (2023), to provide more precise rules than simply summarizing high-reward strategies.
@@ 3.2 Rule-Guided Policy Update (update)
Given a rule $\mathcal{L}_{i}$, the update step produces a new policy $\pi_{i+1}\leftarrow\texttt{update}(\pi_{i},D_{i},\mathcal{L}_{i})$ that is better aligned with $\mathcal{L}_{i}$. There exist many methods for leveraging language instructions to update agents’ policies, though these methods traditionally focus on instructions provided by human experts. Crucially, which method to implement PLLB with depends on the underlying agent and action space representation, and will improve with advances in multi-modal modeling.
@@
For RL policies, we leverage InstructRL (Hu and Sadigh, 2023), which regularizes the learned policy with another policy induced by the linguistic rule $\pi_{\mathcal{L}}$. For instance, in the maze example shown in Figure 1, if the rule is I should go right on every blue cell, the induced policy $\pi_{\mathcal{L}}$ should assign probability 1 to the right action in blue cells, but equal probabilities to all actions in every other situations. In the Q-learning algorithm, this approach simply adds a regularizing term (orange) to the standard Q-learning update rule  [^2] This regularizing term can also be applied in the function approximation setting (Mnih et al., 2013).:
@@
$Q^{\theta}(s_{t},a_{t})\leftarrow r_{\text{t+1}}+\gamma Q^{\theta}(s_{t+1},a_{t+1})~~~\textrm{ where}~~~a_{t+1}=\operatorname*{arg\,max}_{a}~[Q(s_{\text{t+1}},a)~{\color[rgb]{1,.5,0}+~\lambda\log\pi_{\mathcal{L}}(a\mid s_{t})}].$
@@
Here $\gamma$ is a discount factor and $\lambda$ controls the strength of the rule-induced regularization and could be made time-dependent with a pre-defined or learned schedule (e.g., stronger regularization early).
@@
But how do we induce $\pi_{\mathcal{L}}$ from $\mathcal{L}$? Since $\mathcal{L}$ is expressed in natural language, we may obtain a rule-conditional policy by prompting an LM. In particular, our experiments condition the LM on both the current rule and the current state of the agent $s_{t}$, and instruct the LM to generate the next action to obtain a probability distribution over admissible actions (e.g directions), as in Hu and Sadigh (2023). While these prompted LMs may perform tasks poorly on their own (e.g. because of their inability to perform long-range planning or process complex visual input), the regularizer may nonetheless guide Q-learning in the right direction. Running Q-learning updates with the regularization term from experience data gives us a new Q-table $Q^{\theta}_{i+1}$ from which we can derive the new policy $\pi_{i+1}$ by taking actions with maximum expected value in every state $\pi_{i+1}(s_{t})=\text{max}_{a}Q^{\theta}_{t+1}(s_{t},a)$.
@@
Finally, there exist some domains where the policy can be directly implemented by an LM or VLM, such as text-based games explored by methods like ReAct (Yao et al., 2022) or Rememberer (Zhang et al., 2023). In these cases, update can be implemented by conditioning the policy on the rule $\mathcal{L}$ and adding it to the prompt $\pi_{\mathcal{L}}$, ultimately steering the agent’s: $\pi_{i+1}\leftarrow\pi_{\mathcal{L}_{i}}$. Given a new policy $\pi_{i+1}$, we can now close the loop and generate a new rule $\mathcal{L}_{i+1}$.The choice between InstructRL-style regularization and prompt conditioning depends on the action space and policy architecture. In settings with large, continuous action spaces (e.g. reasoning over free-text), conditioning text-generation on rules is likely more tractable, whereas Q-learning style updates may be more appropriate for interactive environments with feedback (e.g. game playing). We implement PLLB with tabular RL agents in Sections 5 and 6, with LM/VLM policies in Section 7, and visuomotor robot controllers in Section 8.
@@ 4 Experiment Set-Up
We run experiments on a diverse set of five tasks, with different agent architectures and action spaces, to showcase how PLLB can train more human-like agents. These include a simple two-player communication game called SaySelect (Section 5), a Maze solving task (Section 6), two collaborative image reconstruction tasks with synthetic (Builder) and natural (Birds) images (Section 7), and grasp planning for a 7-DoF robot (Section 8). For all tasks, we include hyperparameter details (Appendix B.1), the exact prompts used in gen_rule and update (Appendix B.7), and examples of generated rules. For our main experiments, we prioritized using accessible and low-cost open weight language models such as llama-2-70b-chat. However, we analyze the effect of changing the underlying language model on gen_rule in Appendix B.8 and B.9.
To evaluate the cognitive function of PLLB, we first evaluate the policies learned by artificial agents in each environment based on performance metrics such as task reward, similarity to known human-like policies, or generalization. For all tasks, we compare our Bottleneck approach with baseline methods (e.g., standard RL, base LM) and an Adversarial ablation of PLLB that leverages invalid linguistic rules (e.g., obtained by labeling high reward examples as low reward) in gen_rule. The purpose of Adversarial is to test whether PLLB gains arise from meaningful rule content rather than merely introducing any linguistic signal.
To evaluate the communicative function of PLLB, we conduct several human subject studies showing that generated rules $\mathcal{L}$ can help improve human task performance on a variety of metrics (e.g., reward, solving time). Participants were either recruited in-person (for Maze) or on Prolific (Builder, Birds), compensated at a rate of US\$12/hour. All studies were approved by our institution’s IRB. To the best of our knowledge, our work is the first to validate self-generated agent rules from LMs for training human participants.
Finally, we conduct additional sensitivity analysis on the affects of variations in prompt, model temperature, and non-contrastive episode selection in the Appendix.